### v2

# Primordial Power Spectrum of Five Dimensional Uniform Inflation

Abstract
Five dimensional (5D) uniform inflation describes a de Sitter (or approximate) solution of 5D Einstein equations, with cosmological constant and a 5D Planck scale $M_{*}\sim 10^{9}~{}{\rm GeV}$. During the inflationary period all dimensions (compact and non-compact) expand exponentially in terms of the 5D proper time. This set-up requires about 40 $e$-folds to expand the fifth dimension from the fundamental length to the micron size. At the end of 5D inflation (or at any given moment during the inflationary phase) one can interpret the solution in terms of 4D fields using 4D Planck units from the relation $M_{p}^{2}=2\pi RM_{*}^{3}$, which amounts going to the 4D Einstein frame. This implies that if the compactification length $R$ expands $N$ $e$-folds, then the 3D space would expand $23N/$ 2 $e$-folds as a result of a uniform 5D inflation. We reexamine the primordial power spectrum predicted by this model and show that it is consistent with Planck’s measurements of the comic microwave background. The best-fit to Planck data corresponds to $R\sim 10~{}\mu{\rm m}$. A departure of the angular power spectrum predicted by 4D cosmology is expected at multipole moment $\ell\sim 7$.
Recently, we introduced the idea that a compact extra dimension can obtain a large size by five dimensional (5D) uniform inflation, relating the weakness of the actual gravitational force to the size of the observable universe Anchordoqui:2022svl . The requirement of (approximate) flat power spectrum of primordial density fluctuations consistent with observations of the cosmic microwave background (CMB) makes this simple idea possible only for one extra dimension at around the micron scale Anchordoqui:2023etp . Thus, this idea can be naturally combined with the dark dimension proposal for the cosmological constant using the distance/duality conjecture within the swampland program Montero:2022prj . For distances smaller than the compactification length, 5D uniform inflation leads to a scale-invariant Harrison-Zel’dovich spectrum Harrison:1969fb ; Zeldovich:1972zz , because the 2-point function of a massless minimally coupled scalar field (such as a slow-rolling inflaton) in de Sitter space behaves logarithmically at distances larger than the cosmological horizon Ratra:1984yq .
For distances larger than the compactification length (or large angles), the 5D model predicts more power spectrum than standard 4D inflation, corresponding to a nearly vanishing spectral index Anchordoqui:2023etp . Very recently, an investigation was carried out to ascertain the ability of future CMB experiments to constrain or detect cosmological models that modify the CMB power spectra at large angular scales predicted by concordance 4D cosmology Petretti:2024mjy . In this Letter we reinterpret the likelihood analysis carried out in Petretti:2024mjy and show that predictions of 5D uniform inflation are consistent with CMB measurements by the Planck mission Planck:2018nkj . The best-fit to Planck data corresponds to a compactification length $\sim 10~{}\mu{\rm m}$. A departure of the angular power spectrum predicted by 4D cosmology is visible at multipole moment $\ell\sim 7$.
Before proceeding, we pause to note that cosmic variance, which is large at small multipoles (typically below an $\ell$ of order 10), limits the precision of CMB power spectrum measurements. This is due to the finite number of independent measurements that can be used to characterise the temperature variance on large angular scales. These uncertainties make it challenging to define the slope of the power spectrum of temperature fluctuations at low $\ell$. WMAP WMAP:2012fli and Planck Planck:2018nkj data agree with a general up-turn at low $\ell\lesssim 10$, but this is not necessarily a significant result, since all experiments are measuring the same universe.
It is well-known that the power spectrum of the scalar curvature fluctuations predicted by 4D single field inflationary models can be expressed on super-Hubble scales as
$P_{s}(k)=\frac{1}{2M_{p}^{2}\ \varepsilon}\left(\frac{H}{2\pi}\right)^{2}\left(\frac{k}{aH}\right)^{2\delta-\varepsilon}\,,$ (1)
where $k$ is the comoving momentum, $a$ is the cosmic scale factor, $M_{p}=2.48\times 10^{18}~{}{\rm GeV}$ is the reduced Planck mass, $\varepsilon$ and $\delta$ are respectively the first and second slow-roll parameters, and where $H$ characterizes the de Sitter epoch Riotto:2002yw . Note that $H$ and the slow-roll parameters are not constant, but they actually depend very slowly on time. An expansion of $\varepsilon$ and $H$ in terms of the conformal time up to leading order in Hubble flow parameters leads to
$P_{s}(k)=\frac{1}{2M_{p}^{2}\ \varepsilon_{\circledast}}\left(\frac{H_{\circledast}}{2\pi}\right)^{2}\left(\frac{k}{H_{\circledast}}\right)^{2\delta-\varepsilon}\,,$ (2)
where $H_{\circledast}$ and $\varepsilon_{\circledast}$ are the values of the Hubble parameter and Hubble flow function at the expansion’s reference point; for details see Appendix I. Bearing this in mind, the spectrum of CMB anisotropies can be conveniently parametrized by
$P_{s}(k)=A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1}\,,$ (3)
where $A_{s}\simeq 2\times 10^{-9}$ is the scalar amplitude, $n_{s}\simeq 0.96$ is the scalar spectral tilt, and $k_{*}=0.05~{}{\rm Mpc}^{-1}$ is the pivot scale that exits the horizon at $N_{*}\equiv N_{\rm end}-50$ e-folds from the start of inflation ($N_{\rm start}\equiv 0$Planck:2018nkj . [^1] A point worth noting at this juncture is that any given present-day value “$x$” of a CMB scale $k$, including the pivot scale $k_{*}$, should be expressed in the form of $k=a_{\rm today}\,x~{}{\rm Mpc}^{-1}$. However, it is generally assumed that $a_{\rm today}=1$ implicitly, and the scale is expressed as $k=x~{}{\rm Mpc}^{-1}$ Mishra:2021wkm .
Now, the primordial power spectrum for 5D uniform inflation (in Planck units) is found to be
${\cal P}_{s}(k)$ $=$ $\frac{2R_{0}H^{3}}{3\pi^{3}\varepsilon}\Bigg{[}\left(\frac{k}{\hat{a}H}\right)^{2\delta-5\varepsilon}S_{2}(x)+\frac{\varepsilon}{3}\left(\frac{k}{\hat{a}H}\right)^{-3\varepsilon}$ (4)
$\times$ $x^{2}S_{4}(x)\Bigg{]}\,,$
where $\hat{a}$ is the 5D scale factor (normalized such that $\hat{a}_{\rm start}=1$), $R_{0}\sim M_{*}^{-1}$ is the fundamental length scale (coincident with the length of the compact dimension) at the beginning of inflation,
$S_{2}(x)={\rm coth}\ x+x\ {\rm csch}^{2}x\,,$ (5)
and
$S_{4}(x)$ $=$ $15\ {\rm coth}\ x+(4x^{2}\ {\rm coth}^{2}x+12x\ {\rm coth}\ x+15)$ (6)
$\times$ $x\ {\rm csch}^{2}x+2x^{3}\ {\rm csch}^{4}x\,,$
with
$x=\pi kR_{0}=\frac{\pi kR}{\hat{a}_{\rm end}}=\frac{\pi k\widetilde{R}}{a_{\rm end}}=\pi k\ e^{-N}\ \widetilde{R}\,,$ (7)
where $\hat{a}_{\rm end}$ is the 5D scale factor at the end of inflation, $R=e^{2N/3}R_{0}$ is the value of the compactification length at the end of inflation (around micron), $\tilde{R}=e^{N}R_{0}$ is the “corrected compactification scale” on the brane (around kilometer), and $N$ is the number of $e$-folds in 4D Antoniadis:2023sya .
Duplicating the expansion procedure on $H$ and $\varepsilon$ we obtain
${\cal P}_{s}(k)=\frac{2R_{0}H_{\circledast}^{3}}{3\pi^{3}\varepsilon_{\circledast}}\ \left(\frac{k}{H_{\circledast}}\right)^{2\delta-5\varepsilon}S_{2}(x),$ (8)
where we have neglected the second term in (4) because it is largely suppressed by the slow-roll parameter when compared to the first term.
In the limit $kR_{0}\gg 1$ the spectrum (8) can be recast as
${\cal P}_{s}(k)\underset{R_{0}k\gg 1}{\simeq}\frac{R_{0}H_{\circledast}^{3}}{3\pi^{2}\varepsilon_{\circledast}}\biggl{(}\frac{k}{H_{\circledast}}\biggr{)}^{2\delta-5\varepsilon}\,,$ (9)
whereas for $kR_{0}\ll 1$, the spectrum (8) can be rewritten as
${\cal P}_{s}(k)\underset{R_{0}k\ll 1}{\simeq}\frac{2H_{\circledast}^{3}}{3\pi^{3}\varepsilon_{\circledast}k}\ \biggl{(}\frac{k}{H_{\circledast}}\biggr{)}^{2\delta-5\varepsilon}\,.$ (10)
To a first approximation we can also ignore the pre-factor boosting the scalar amplitude because the slow roll parameters are very small and so the asymptotic expression (9) gives a scale invariant spectrum
${\cal P}_{s_{\gg}}(k)\sim{\cal A}_{s_{\gg}}\,,$ (11)
while (10) leads to
${\cal P}_{s_{\ll}}(k)\sim\frac{1}{k}\ {\cal A}_{s_{\ll}}\,,$ (12)
where ${\cal A}_{s_{\gg}}$ and ${\cal A}_{s_{\gg}}$ are the corresponding amplitudes of the two asymptotic regimes.
Table 1: Priors on the cosmological parameters.
         Cosmological Parameter          Priors
$\Omega_{b}h^{2}$ [0.02, 0.0265]0.020.0265[0.02,\,0.0265][ 0.02 , 0.0265 ]
$\Omega_{c}h^{2}$ [0.1, 0.135]0.10.135[0.1,\,0.135][ 0.1 , 0.135 ]
$100\theta_{s}$ [1.03, 1.05]1.031.05[1.03,\,1.05][ 1.03 , 1.05 ]
$\tau_{\rm reio}$ [0.03, 0.08]0.030.08[0.03,\,0.08][ 0.03 , 0.08 ]
$n_{s}$ [0.920, 0.996]0.9200.996[0.920,\,0.996][ 0.920 , 0.996 ]
$\ln(10^{10}A_{s})$ [2.763, 4.375]2.7634.375[2.763,\,4.375][ 2.763 , 4.375 ]
$\ln(10^{10}{\cal A}_{s})$ [2.763, 4.375]2.7634.375[2.763,\,4.375][ 2.763 , 4.375 ]
$\zeta$ $[-6.6,\,-2.1]$
In Anchordoqui:2023etp we adopted a $\Theta$-function approximation to match the two asymptotic expressions and showed that the resulting spectrum is partially consistent with CMB data, including effects of cosmic variance at large angles. However, the use of a step $\Theta$-function approximation to accommodate the change of behaviour amounts to throwing away any memory at smaller angles and just parametrize the low-multipole region with an approximate $1/k$ fit.
Next, in line with our stated plan, we reinterpret the results of the likelihood analysis carried out in Petretti:2024mjy using Planck 2018 CMB data on temperature and $E$-mode polarization Planck:2018nkj . This analysis compares predictions from the widely accepted spatially-flat $\Lambda$ cold dark matter (CDM) model (supplemented by an initial stage of slow-roll inflation) and 5D uniform inflation.
$\Lambda$CDM requires only 6 independent parameters,
$\mathscr{P}_{4{\rm D}}=\{\Omega_{b}h^{2},\,\Omega_{c}h^{2},\,\theta_{s},\,\tau_{\rm reio},\,n_{s},\,A_{s}\}\,,$ (13)
to completely specify the cosmological evolution, where $\Omega_{b}$ is the baryon density, $\Omega_{c}$ is the CDM density, $\theta_{s}$ is the angular size of the sound horizon at recombination, $\tau_{\rm reio}$ is the Thomson scattering optical depth due to reionization, $n_{s}$ is the scalar spectral tilt, $A_{s}$ is the power spectrum amplitude of adiabatic scalar perturbations, and $h=H_{0}/(100~{}{\rm km}\;{\rm s^{-1}}{\rm Mpc}^{-1})$ is the dimensionless Hubble constant. The $\Omega_{i}$ parameters are defined as the ratio of the present day mean density of each component $i$ to the critical density.
Figure 1: Primordial power spectra of the best-fit candidates from the likelihood analysis of Ref. Petretti:2024mjy . The featureless best-fit (consistent with the concordance model of cosmology) is represented by the black dashed line and the best-fit of 5D uniform inflation by the solid line.
The primordial power spectrum predicted by 5D uniform inflation given in (8) has been normalized such that $a_{\rm start}=1$. To accomodate the normalization of CMB data with $a_{\rm today}=1$ we first need to rescale the physical size of the causal patch at the beginning of inflation $R_{0}$ to today ${\cal R}_{\rm today}$. After the proper rescaling of $R_{0}$ has been worked out, the primordial power spectrum of 5D uniform inflation can be parametrized by
${\cal P}_{s}(k)={\cal A}_{s}\ \left(\frac{k}{k_{*}}\right)^{n_{s}-1}\ S_{2}(\pi k{\cal R}_{\rm today})\,,$ (14)
where
${\cal R}_{\rm today}=\frac{\tilde{R}}{a_{\rm end}}=\frac{1}{\pi k_{*}e^{\zeta}}\,,$ (15)
with
$a_{\rm end}=\left(\frac{M_{I}}{10^{9}~{}{\rm GeV}}\right)^{-1}2.3\times 10^{-22}$ (16)
the scale factor at the end of inflation (for $a_{\rm today}=1$), $M_{I}$ the inflation scale, and $\zeta$ a free parameter of the model to be determined by fitting the data. The 5D setup then requires 7 parameters,
$\mathscr{P}_{5{\rm D}}=\{\Omega_{b}h^{2},\,\Omega_{c}h^{2},\,\theta_{s},\,\tau_{\rm reio},\,n_{s},\,A_{s},\,\zeta\}\,,$ (17)
to specify the cosmological evolution.
The analysis of Petretti:2024mjy relies on uniform priors on the cosmological parameters, which are listed in Table 1. Note that the priors on the $\zeta$ parameter guarantee that $S(\pi k_{*}{\cal R}_{\rm today})=1$, and so ${\cal P}_{s}(k_{*})={\cal A}_{s}$. The best-fit to the data gives $\zeta=-6.44$. Substituting the best-fit value of $\zeta$ into (15) we obtain ${\cal R}_{\rm today}\sim 4~{}{\rm Gpc}$. Now, substituting (16) into (15) with $M_{I}\sim M_{*}\sim 10^{9}~{}{\rm GeV}$ and ${\cal R}_{\rm today}\sim 4~{}{\rm Gpc}$ leads to $\tilde{R}\sim 27~{}{\rm km}$. Converting $\tilde{R}$ to higher-dimensional units it is scaled down by an additional factor $M_{*}/M_{p}$, which implies that the compactification length at the end of inflation is $R\sim 10~{}\mu{\rm m}$.
To determine whether there is an improvement in the likelihood of 5D uniform inflation over 4D cosmology, the analysis of Petretti:2024mjy reports the Bayes factor defined as the ratio of evidences between two models. The Bayes factor is calculated as
$\ln B=\ln{\cal Z}_{5{\rm D}}-\ln{\cal Z}_{4{\rm D}}=-1.45\pm 0.35\,,$ (18)
which implies that the 5D model is statistically slightly disfavored compared to the standard 4D scenario. Besides, values of $\zeta>-5.38$ are ruled out by the data at 95%CL. This leads to the following 95%CL lower limits: ${\cal R}_{\rm today}>1.4~{}{\rm Gpc}$ and $R>4~{}\mu{\rm m}$.
In Fig. 1 we show a comparison between the primordial power spectra of the concordance 4D model of cosmology (dashed line) and 5D uniform inflation (solid line) from the likelihood analysis of Ref. Petretti:2024mjy . One can check by inspection that the primordial power spectrum predicted by 5D uniform inflation departs from the 4D $\Lambda$CDM prediction at around $k\sim 2/{\cal R}_{\rm today}\sim 5\times 10^{-4}~{}{\rm Mpc}^{-1}$. This corresponds to a multipole moment $\ell\simeq kd_{A}\simeq 7$, where $d_{A}=1.4\times 10^{4}~{}{\rm Mpc}$ is the angular diameter distance to the last scattering surface Bridle:2003sa . An explicit comparison between the different interpretations of the result from the likelihood analysis is provided in Appendix II.
Lastly, it is constructive to connect with contrasting and complementary perspectives to comment on two caveats of the study presented herein:
  • When $k$ exits the 5D horizon $R$ does not have the micron-size yet. We have assumed that this fact does not significantly modify the predictions of (8).
  • The 5D horizon is different from the 4D horizon and this may introduce an additional effect/time-dependence, which we have assumed can be neglected.
A deeper investigation along these lines is obviously important to be done.
In summary, we have reinterpreted the likelihood analysis carried out in Petretti:2024mjy and showed that predictions of 5D uniform inflation are consistent with CMB measurements by the Planck mission. In particular, the best-fit to the data corresponds to $R\sim 10~{}\mu{\rm m}$. A departure of the angular power spectrum predicted by 4D cosmology is visible at multipole moment $\ell\sim 7$. Future data from LiteBIRD LiteBIRD:2022cnt and CMBS-4 CMB-S4:2016ple will provide a decisive test for the ideas discussed in this Letter.
Note Added
Within the dark dimension scenario cosmological overproduction of bulk graviton modes leads to a normalcy temperature $T_{*}\sim 1~{}{\rm GeV}$ Arkani-Hamed:1998sfv ; Gonzalo:2022jac . The behavior of the scale factor between the end of inflation and $T_{*}$ is model dependent. In (16) we adopted a simple radiation-like behavior for the scale factor between $M_{I}$ and $T_{*}$, but of course assuming that the universe must be free of bulk modes above $T_{*}$. As noted in Macesanu:2004gf , one possible pathway to accommodate our assumption could be the softening of the gravity-matter interaction due to brane fluctuations Bando:1999di ; Bando:2000ch . Indeed, if brane oscillations were to be excited above the normalcy temperature, then the matter-graviton interaction vertex would aquire an effective form factor ${\cal F}=e^{-\frac{1}{2}\frac{m^{2}}{\Delta^{2}}}$, where $m$ is the mass of the Kaluza-Klein graviton and $\Delta\sim 4\pi\sqrt{\tau}/M_{*}$ is the “softening scale”, which is related to the brane tension $\tau$ Murayama:2001av . As a consequence, the cross-section for production of gravitons could be exponentially suppressed with a suitably small tension of our world, i.e., for a relatively soft brane. It is interesting to note that the cooling process of Supernova 1987A leads to the constraint $\tau\gtrsim 100~{}{\rm GeV}^{4}$ Kugo:1999mf . An investigation along this line is obviously important to be done.
Acknowledgements
The work of L.A.A. is supported by the U.S. National Science Foundation (NSF Grant PHY-2412679). I.A. is supported by the Second Century Fund (C2F), Chulalongkorn University.
Appendix I
An expansion of the Hubble flow functions $\varepsilon_{i}$ and Hubble parameter $H$ in terms of the conformal time $\tau=(aH)^{-1}$ around a given time $\tau_{\circledast}$ leads to
$\varepsilon_{i}$ $=$ $\varepsilon_{i}^{\circledast}\Bigg{\{}1-\ln\left(\frac{\tau}{\tau_{\circledast}}\right)\varepsilon_{i+1}^{\circledast}\big{(}1+\varepsilon_{1}^{\circledast}+\varepsilon_{1}^{\circledast 2}+\varepsilon_{1}^{\circledast}\varepsilon_{2}^{\circledast}\big{)}+\frac{1}{2}\left[\ln\left(\frac{\tau}{\tau_{\circledast}}\right)\right]^{2}\varepsilon_{i+1}^{\circledast}\big{(}\varepsilon_{i+1}^{\circledast}+\varepsilon_{i+2}^{\circledast}+\varepsilon_{1}^{\circledast}\varepsilon_{2}^{\circledast}+2\varepsilon_{1}^{\circledast}\varepsilon_{i+1}^{\circledast}+2\varepsilon_{1}^{\circledast}\varepsilon_{i+2}^{\circledast}\big{)}$ (19)
$\qquad~{}-\frac{1}{6}\left[\ln\left(\frac{\tau}{\tau_{\circledast}}\right)\right]^{3}\varepsilon_{i+1}^{\circledast}\big{(}\varepsilon_{i+1}^{\circledast 2}+3\varepsilon_{i+1}^{\circledast}\varepsilon_{i+2}^{\circledast}+\varepsilon_{i+2}^{\circledast 2}+\varepsilon_{i+2}^{\circledast}\varepsilon_{i+3}^{\circledast}\big{)}\Bigg{\}}+\mathcal{O}\big{(}\varepsilon^{\circledast 5}\big{)}$
and
$H$ $=$ $H_{\circledast}\Bigg{\{}1+\ln\left(\frac{\tau}{\tau_{\circledast}}\right)\varepsilon_{1}^{\circledast}\big{(}1+\varepsilon_{1}^{\circledast}+\varepsilon_{1}^{\circledast 2}+\varepsilon_{1}^{\circledast}\varepsilon_{2}^{\circledast}\big{)}+\frac{1}{2}\left[\ln\left(\frac{\tau}{\tau_{\circledast}}\right)\right]^{2}\varepsilon_{1}^{\circledast}\big{(}\varepsilon_{1}^{\circledast}-\varepsilon_{2}^{\circledast}+2\varepsilon_{1}^{\circledast 2}-3\varepsilon_{1}^{\circledast}\varepsilon_{2}^{\circledast}\big{)}$ (20)
$\qquad~{}+\frac{1}{6}\left[\ln\left(\frac{\tau}{\tau_{\circledast}}\right)\right]^{3}\varepsilon_{1}^{\circledast}\big{(}\varepsilon_{1}^{\circledast 2}-3\varepsilon_{1}^{\circledast}\varepsilon_{2}^{\circledast}+\varepsilon_{2}^{\circledast 2}+\varepsilon_{2}^{\circledast}\varepsilon_{3}^{\circledast}\big{)}\Bigg{\}}+\mathcal{O}\big{(}\varepsilon^{\circledast 4}\big{)}\,,$
where $\varepsilon=\varepsilon_{1}$ and $\varepsilon_{2}=-2\delta+2\varepsilon$ Antoniadis:2024abm . Using (19) and (20) at leading order it is easily seen that
$\frac{H^{2}}{\varepsilon}\left(\frac{k}{aH}\right)^{2\delta-\varepsilon}\sim\frac{H_{\circledast}^{2}}{\varepsilon_{\circledast}}\left(\frac{k}{H_{\circledast}}\right)^{2\delta-\varepsilon}$ (21)
and that
$\frac{H^{3}}{\varepsilon}\left(\frac{k}{\hat{a}H}\right)^{2\delta-5\varepsilon}\sim\frac{H_{\circledast}^{3}}{\varepsilon_{\circledast}}\left(\frac{k}{H_{\circledast}}\right)^{2\delta-5\varepsilon}\,.$ (22)
Appendix II
The discussion in this Appendix provides a critical assessment of the likelihood analysis presented in the first version of Petretti:2024mjy , submitted to the arXiv on November 2024.
In Petretti:2024mjy the authors first establish the relation
$R_{0}=\frac{1}{\pi k_{*}e^{\xi}}\,,$ (23)
and then claim that since the length of the compact dimension at the beginning of inflation $R_{0}$ is unknown, it can be determined through a likelihood fit to the CMB data. In this Appendix we show that this procedure leads to an ambiguity, and that actually $R_{0}$ cannot be inferred from the fit.
It is well-known that in ordinary 4D inflation, where there is no extra scale accounting for the evolution of the compact space, the power spectrum depends on the spectral index and this gives the power of $k/k_{*}$ for any reference momentum $k_{*}$. In other words, $k_{*}$ is arbitrary. Likewise, after the end of 5D uniform inflation the compactification length remains fixed on the micron scale and therefore a CMB data analysis can be considered effectively 4D. However, for 5D uniform inflation, an ambiguity emerges when backtracking the expansion of the universe beyond the end of inflation. To understand why this is the case, we recall that we know how $R$ grows in time from $R_{0}$ during inflation, but it is important to stress that the comoving momentum $k$ is an extra variable. Now, the relation between the times each $k$ exits the horizon and the size of the compact space $R$ at that time is known, but because $k$ is an extra variable we do not know at which time a given $k$ would exit the horizon.
A crucial step in defining (7) and subsequently (15) is that the wavelength fluctuation $k$ exits the horizon precisely at the end of inflation. This allows us to relate the free parameter in the fit with ${\cal R}_{\rm today}$ via (15). We reiterate that for the best fit value $\xi=-6.44$, yielding ${\cal R}_{\rm today}\sim 4~{}{\rm Gpc}$. Note that from the illogical relation (23) it follows that $R_{0}\sim 4~{}{\rm Gpc}$. Clearly, the size of the compact space $R_{0}\sim M_{*}^{-1}$ cannot be of Gpc scale at the beginning of inflation. We conclude that the interpretation of the result from the fit in Petretti:2024mjy using (23) is misleading. On the other hand, assuming that the fit in Petretti:2024mjy is correct, we reinterpreted the result using (15), for which ${\cal R}_{\rm today}$ is of order ${\rm Gpc}$, leading to a characteristic distance at the CMB epoch of ${\cal O}({\rm Mpc})$ and to a compact space at the end of inflation of $\sim 10~{}\mu{\rm m}$. This corresponds to an angular scale of $\sim 10^{\circ}$ in the sky, for which the uncertainties in the angular power spectrum become large.